OpenStudy (anonymous):

What is the distance between –4 and 3?

OpenStudy (saifoo.khan):

7

hero (hero):

lame

hero (hero):

No steps, huh?

hero (hero):

Same ol' Saifoo

OpenStudy (anonymous):

yeah as ther's only one variable co-ordinate is given so saifoo is right

hero (hero):

I know it's right. That's not what my complaint is about

OpenStudy (anonymous):

3-(-4) = 3+4 = 7